# Understanding World or Predicting Future? A Comprehensive Survey of World Models

### v2

@@ Abstract.
The concept of world models has garnered significant attention due to advancements in multimodal large language models such as GPT-4 and video generation models such as Sora, which are central to the pursuit of artificial general intelligence. This survey offers a comprehensive review of the literature on world models. Generally, world models are regarded as tools for either understanding the present state of the world or predicting its future dynamics. This review presents a systematic categorization of world models, emphasizing two primary functions: (1) constructing internal representations to understand the mechanisms of the world, and (2) predicting future states to simulate and guide decision-making. Initially, we examine the current progress in these two categories. We then explore the application of world models in key domains, including autonomous driving, robotics, and social simulacra, with a focus on how each domain utilizes these aspects. Finally, we outline key challenges and provide insights into potential future research directions. We summarize the representative papers along with their code repositories in https://github.com/tsinghua-fib-lab/World-Model.
World model, model-based RL, video generation, embodied environment, autonomous driving, robots, social simulacra
*These two authors contributed equally.
@@
copyright: acmcopyrightjournalyear: 20xxjournal: CSURcopyright: noneccs: Computing methodologies Machine learningccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Modeling and simulation
@@ 1. Introduction
The main contributions of this survey can be summarized as follows: (1) We present a novel categorization system for world models structured around two primary functions: constructing implicit representations to understand the mechanism of the external world and predicting future states of the external world. The first category focuses on the development of models that learn and internalize world knowledge to support subsequent decision-making, while the latter emphasizes enhancing predictive and simulative capabilities in the physical world from visual perceptions. (2) Based on this categorization, we classify how various key application areas, including autonomous driving, robots, and social simulacra, emphasize different aspects of world models. (3) We highlight future research directions and trends of world models that can adapt to a broader spectrum of practical applications.
@@ 2. Background and Categorization
The concept of building an internal model of the world has a long history in AI, dating back to foundational work such as Marvin Minsky’s frame representation in the 1960s (minsky1974framework, ), designed to systematically capture structured knowledge about the world. Ha et al. (ha2018recurrent, ; ha2018world, ) significantly revived and popularized the term “world model” in 2018 by proposing neural-network-based implicit models for learning latent representations. This line of research aligns with the psychological theory of “mental models” (johnson1983mental, ) [^1] https://plato.stanford.edu/entries/mental-representation/, which holds that humans perceive the external world by abstracting it into simplified elements and relationships—an underlying philosophical root reflected alike in both frames and world models. This principle suggests that our descriptions of the world, when viewed from a deep, internal perspective, typically involve constructing an abstract representation that suffices without requiring detailed depiction. Building upon this conceptual framework, the authors introduce an agent model inspired by the human cognitive system, as illustrated in Figure 1. In this pioneering model, the agent receives feedback from the real-world environment, which is then transformed into a series of inputs that train the model. This model is adept at simulating potential outcomes following specific actions within the external environment. Essentially, it creates a mental simulation of potential future world evolutions, with decisions made based on the predicted outcomes of these states. This methodology closely mirrors the Model-based Reinforcement Learning (MBRL) method, where both strategies involve the model generating internal representations of the external world. These representations facilitate navigation through and resolution of various decision-making tasks in the real world.
@@
In the visionary article on the development of autonomous machine intelligence in 2022 (lecun2022path, ), Yann LeCun introduced the Joint Embedding Predictive Architecture (JEPA), a framework mirroring the human brain’s structure. As illustrated in Figure 1, JEPA comprises a perception module that processes sensory data, followed by a cognitive module that evaluates this information, effectively embodying the world model. This model allows the brain to assess actions and determine the most suitable responses for real-world applications. LeCun’s framework is intriguing due to its incorporation of the dual-system concept, mirroring ”fast” and ”slow” thinking. System 1 involves intuitive, instinctive reactions: quick decisions made without a world model, such as instinctively dodging an oncoming person. In contrast, System 2 employs deliberate, calculated reasoning that considers the future state of the world. It extends beyond immediate sensory input, simulating potential future scenarios, like predicting events in a room over the next ten minutes and adjusting actions accordingly. This level of foresight requires constructing a world model to effectively guide decisions based on the anticipated dynamics and evolution of the environment. In this framework, the world model is essential for understanding and representing the external world. It models the state of the world using latent variables, which capture key information while filtering out redundancies. This approach allows for a highly efficient, minimalistic representation of the world, facilitating optimal decision-making and planning for future scenarios.
The ability of models to capture world knowledge is critical for their effective performance in a wide range of real-world tasks. In the recent wave of works on large language models starting from 2023, several have demonstrated the presence of latent world knowledge. In other words, these models capture intuitive knowledge, including spatial and temporal understanding, which enables them to make predictions about real-world scenarios (gurnee2023language, ; manvi2023geollm, ). Furthermore, LLMs are capable of modeling the external world through cognitive maps, as indicated by recent research revealing the brain-like structures embedded within them (li2024geometryconceptssparseautoencoder, ). These models can even learn to predict future events based on prior experiences, thereby enhancing their utility and applicability in real-world contexts.
The above world models primarily represent an implicit understanding of the external world. Powered by generative learning like diffusion modeling and model architecture like transformer, recent video generation models (e.g., Sora (sora2024, ), Keling (keling2024, ), Gen-2 (cui2024feasibility, ), etc.) take text instructions or real-world visual data as input and output high-quality video frames. Notably, these models demonstrate exceptional modeling capabilities, such as maintaining consistency in 3D video simulations, producing physically plausible outcomes, and simulating digital environments. These capabilities suggest that they not only mimic the appearance of but also model the real-world dynamics within simulation scenarios, focusing on realistically modeling dynamic world changes rather than merely representing static world states.
@@
  • Applications of world models (Section 5): World models have a wide range of applications across various fields, including autonomous driving, robotics, and social simulacra. We explore how the integration of world models in these domains advances both theoretical research and practical implementations, emphasizing their transformative potential in real-world applications.
@@ 3.1.1. World model in model-based RL
(1) $\min_{\theta}\mathbb{E}_{s^{\prime}\sim M^{*}(\cdot|s,a)}[||s^{\prime}-M_{\theta}(s,a)||^{2}_{2}],$
@@ 3.2. World Knowledge Learned by Models
GeoLLM (manvi2023geollm, ) 2024 ICLR Language Learning
GeoLLM-Bias (manvi2024large, ) 2024 ICML Language Learning
@@
CityGPT (feng2024citygpt, ) 2024 arxiv Language Learning
CityBench (feng2024citybench, ) 2024 arxiv Language&Vision Benchmark
@@
After pretraining on large-scale web text and books (touvron2023llama, ; chatgpt, ), large language models attain extensive knowledge about the real world and common sense relevant to daily life. This embedded knowledge is considered crucial for their remarkable ability to generalize and perform effectively in real-world tasks. For instance, researchers leverage the common sense of large language models for task planning (zhao2024large, ), robot control (huang2022inner, ), and image understanding (liu2024visual, ). Furthermore, Li et al.(li2024geometryconceptssparseautoencoder, ) discover brain-like structures of world knowledge embedded in the high-dimensional vectors that represent the universe of concepts in large language models. Also, Li et al.(li2024vision, ) demonstrate that language models partially converge towards representations isomorphic to those of vision models.
@@ 3.2.1. Knowledge of the Global Physical World
While large language models do acquire some implicit knowledge of the real world (gurnee2023language, ; li2024geometryconceptssparseautoencoder, ), the quality of this knowledge remains questionable (roberts2023gpt4geo, ; feng2024citygpt, ). For example, Feng et al. (feng2024citygpt, ) find that the urban knowledge embedded in large language models is often coarse and inaccurate. To address this, they propose an effective framework to improve the acquisition of urban knowledge of specific cities in large language models. From these works, we can see that although large language models have demonstrated the ability to capture certain aspects of real-world knowledge (gurnee2023language, ; li2024geometryconceptssparseautoencoder, ; roberts2023gpt4geo, ), it is clear that further efforts are needed to enhance this knowledge to enable broader and more reliable real-world applications (feng2025survey, ).
@@ 3.2.2. Knowledge of the Local Physical World
Recent studies explore actively encouraging models to learn abstract knowledge through cognitive map-like processes across various environments. For example, Cornet et al. (gornet2024automated, ) show that in a simplified Minecraft world, visual predictive coding lets an agent build a spatial cognitive map purely from pixels. Once trained, the latent map encodes its metric distance to any target, enabling accurate rollout of future observations. Lin et al. (lin2024learningmodelworldlanguage, ) investigate teaching models to understand the game environments through a world model learning procedure, specifically by predicting the subsequent frame of the environment. In this way, the model can generate better actions in dynamic environments. Moreover, Jin et al. (jinemergent, ) find that language models can learn the emergent representations of program semantics by predicting the next token.

### v4

@@ Abstract.
The concept of world models has garnered significant attention due to advancements in multimodal large language models such as GPT-4 and video generation models such as Sora, which are central to the pursuit of artificial general intelligence. This survey offers a comprehensive review of the literature on world models. Generally, world models are regarded as tools for either understanding the present state of the world or predicting its future dynamics. This review presents a systematic categorization of world models, emphasizing two primary functions: (1) constructing internal representations to understand the mechanisms of the world, and (2) predicting future states to simulate and guide decision-making. Initially, we examine the current progress in these two categories. We then explore the application of world models in key domains, including generative games, autonomous driving, robotics, and social simulacra, with a focus on how each domain utilizes these aspects. Finally, we outline key challenges and provide insights into potential future research directions. We summarize the representative papers along with their code repositories in https://github.com/tsinghua-fib-lab/World-Model.
*These authors contributed equally.
@@
copyright: acmcopyrightjournalyear: 20xxjournal: CSURcopyright: none
@@ 1. Introduction
The main contributions of this survey can be summarized as follows: (1) We present a novel categorization system for world models structured around two primary functions: constructing implicit representations to understand the mechanism of the external world and predicting future states of the external world. The first category focuses on the development of models that learn and internalize world knowledge to support subsequent decision-making, while the latter emphasizes enhancing predictive and simulative capabilities in the physical world from visual perceptions. (2) Based on this categorization, we classify how various key application areas, including generative games, autonomous driving, robots, and social simulacra, emphasize different aspects of world models. (3) We highlight future research directions and trends of world models that can adapt to a broader spectrum of practical applications.
@@ 2. Background and Categorization
Figure 1. The roadmap of world models in deep learning era.
2.1. History and Current Development
@@
Pre Deep Learning Era. The concept of building an internal model of the world has a long history in AI, dating back to foundational work such as Marvin Minsky’s frame representation in the 1960s (minsky1974framework, ), designed to systematically capture structured knowledge about the world. In the context of reinforcement learning, world models emerged as a fundamental component of model-based approaches, where agents construct explicit representations of their environment’s dynamics. Early work in this domain focused on learning transition models that could predict the next state given the current state and action (sutton1990integrated, ), enabling agents to perform planning and simulate potential action sequences before execution. These environment models were typically represented using tabular methods or simple parametric functions, laying the groundwork for more sophisticated world modeling approaches that would emerge with the advent of deep learning.
@@
Model-based Reinforcement Learning. Ha et al. (ha2018recurrent, ; ha2018world, ) significantly revived and popularized the term “world model” in 2018 by proposing a recurrent neural-network-based implicit model for learning latent representations. This line of research aligns with the psychological theory of “mental models” (johnson1983mental, ) [^1] https://plato.stanford.edu/entries/mental-representation/, which posits that humans perceive the external world by abstracting it into simplified elements and relationships—an underlying philosophical principle reflected in both frames and world models. This principle suggests that our understanding of the world, when viewed from a cognitive perspective, typically involves constructing abstract representations that capture essential patterns without requiring exhaustive detail. Building upon this conceptual framework, the authors introduce an agent module inspired by the human cognitive system, as illustrated in Figure 2. In this recurrent world model (RWM), the agent receives feedback from the real-world environment, which is then transformed into a series of inputs that train the model. This model is adept at simulating potential outcomes following specific actions within the external environment. Essentially, it creates a mental simulation of potential future world evolutions, with decisions made based on the predicted outcomes of these states. This methodology closely mirrors the model-based reinforcement learning method, where both strategies involve the model generating internal representations of the external world to facilitate navigation through and resolution of various decision-making tasks. Following this conceptual foundation, subsequent developments have further advanced world model architectures, including Google DeepMind’s Dreamer series (hafner2019dream, ; hafner2020mastering, ; hafner2025mastering, ), which has demonstrated the scalability and effectiveness of learned world representations across increasingly complex domains.
@@
Self-supervised Learning. In the visionary article on the development of autonomous machine intelligence in 2022 (lecun2022path, ), Yann LeCun introduced the Joint Embedding Predictive Architecture (JEPA), a framework mirroring the human brain’s structure. As illustrated in Figure 2, JEPA comprises a perception module that processes sensory data (i.e., an encoder), followed by a cognitive module (i.e., a predictor) that evaluates this information, effectively embodying the world model. This model allows the brain to assess actions and determine the most suitable responses for real-world applications. A key innovation of JEPA lies in its self-supervised learning paradigm, which enables the system to learn rich representations of the world without relying on extensive labeled data. Rather than predicting raw sensory inputs in pixel space, JEPA learns to predict abstract representations in a latent embedding space, making the learning process more efficient and robust. This approach allows the model to capture semantic relationships and causal structures in the data while avoiding the computational burden and potential pitfalls of pixel-level prediction, such as focusing on irrelevant details or noise. LeCun’s framework is particularly intriguing due to its incorporation of the dual-system concept, mirroring ”fast” and ”slow” thinking. System 1 involves intuitive, instinctive reactions: quick decisions made without explicit world model consultation, such as instinctively dodging an oncoming person. In contrast, System 2 employs deliberate, calculated reasoning that leverages the learned world model to consider future states. It extends beyond immediate sensory input, simulating potential future scenarios through the self-supervised representations, like predicting events in a room over the next ten minutes and adjusting actions accordingly. This level of foresight requires constructing a world model that can effectively guide decisions based on the anticipated dynamics and evolution of the environment. In this framework, the world model is essential for understanding and representing the external world through self-supervised learning of latent variables, which capture key information while filtering out redundancies. This approach allows for a highly efficient, minimalistic representation of the world, facilitating optimal decision-making and planning for future scenarios. Building upon these principles, recent implementations such as V-JEPA (bardes2024revisiting, ) and V-JEPA2 (assran2025v, ) have demonstrated the practical viability of video-based self-supervised learning, showing how JEPA architectures can learn rich spatiotemporal representations from unlabeled video data for downstream vision tasks.
Large Language Models. ”The limits of my language mean the limits of my world.”—Ludwig Wittgenstein. This profound observation finds particular relevance in the context of large language models, which learn fundamental principles of world operation through textual data that can be leveraged to construct comprehensive world models. Recent research has demonstrated that LLMs trained on vast corpora naturally acquire latent world knowledge, including spatial and temporal understanding, enabling them to make sophisticated predictions about real-world scenarios (gurnee2023language, ; manvi2023geollm, ). This capability has been harnessed for model-based task planning, where pre-trained language models serve as the foundation for constructing world models that can reason about complex sequential tasks (guan2023leveraging, ). The integration of multimodal capabilities further enhances world modeling potential. Multimodal Large Language Models (MLLMs) can process and integrate information across visual, textual, and other sensory modalities, creating richer and more comprehensive world representations (ge2024worldgpt, ). Understanding how these models process, represent, and utilize world knowledge remains crucial for developing more effective world models that can bridge the gap between linguistic knowledge and real-world understanding (yang2025thinking, ).
Video Generation. Video generation has emerged as the predominant approach to world modeling in contemporary AI research. Unlike earlier implicit world representations, these models explicitly generate visual sequences that demonstrate understanding of temporal dynamics, spatial consistency, and physical laws. Powered by advanced generative techniques such as diffusion modeling and transformer architectures, recent video generation models including Sora (sora2024, ), Keling (keling2024, ), and Gen-2 (runway2023, ) take text instructions or real-world visual data as input and produce high-quality video sequences. These models demonstrate exceptional world modeling capabilities, including maintaining consistency in 3D video simulations, producing physically plausible outcomes, and simulating complex digital environments. The sophistication of these approaches suggests they model underlying real-world dynamics rather than merely generating visually appealing content. This represents a fundamental shift toward world models that can actively simulate and predict how environments evolve over time. Recent developments have further advanced this paradigm, with Cosmos (agarwal2025cosmos, ) achieving breakthrough performance in physics law adherence and Genie 3 (genie3, ) enabling real-time interaction capabilities for controllable world simulation.
Interactive 3D Environments. Interactive 3D scene generation represents another important paradigm in world modeling, focusing on creating immersive 3D worlds that enable spatial exploration and user interaction within virtual environments. Representative work such as Wonderworld demonstrates the capability to generate interactive 3D scenes from a single 2D image (yu2025wonderworld, ), showcasing the potential for creating explorable virtual worlds from minimal input. This approach emphasizes spatial consistency, geometric understanding, and real-time responsiveness to user navigation and interaction. Recent advances have significantly expanded these capabilities, with Matrix-3D achieving wide-coverage omnidirectional explorable 3D world generation through panoramic 3D reconstruction (yang2025matrix, ), and HunyuanWorld 1.0 enabling immersive 360° experiences through semantically layered 3D mesh representations that provide seamless compatibility with existing computer graphics pipelines (team2025hunyuanworld, ).
Applications. World models have rapidly expanded across diverse application domains since 2023. In autonomous driving, foundational works such as GAIA-1 and Drive-WM established approaches for modeling vehicle interactions and environmental dynamics in complex traffic scenarios (hu2023gaia, ; wang2023drivingfuturemultiviewvisual, ). The robotics domain has similarly advanced, exemplified by DayDreamer in 2023 (wu2023daydreamer, ) and continuing with recent developments in 2025 for robotic manipulation tasks (lu2025gwm, ). Navigation applications have emerged with robot path planning (bar2025navigation, ) extending to six-degree-of-freedom aerial agents (zhao2025airscape, ). Gaming represents a particularly promising domain, with landmark work on world and human action models (WHAM) demonstrating how world models can create dynamic, responsive virtual environments (kanervisto2025world, ). At the larger scale, agent-based social simulation leverages world models to understand complex societal dynamics and human interactions, offering computational insights into real-world social phenomena (piao2025agentsociety, ).
2.2. Evolving Concept from Multiple Domains
The concept of world models in artificial intelligence has deep psychological roots that extend far beyond contemporary machine learning. Understanding these foundational connections reveals how modern AI world models represent a computational realization of fundamental cognitive principles that have been studied for decades across multiple disciplines.
The psychological concept of mental models was first articulated by Scottish psychologist Kenneth Craik in his seminal work ”The Nature of Explanation” (1943), where he proposed that ”the mind constructs small-scale models of reality” to predict and understand external events (craik1943nature, ). Craik’s insight was that human cognition fundamentally operates by creating internal representations that capture the essential structure and dynamics of the external world, enabling predictive reasoning and adaptive behavior.
This foundational concept was systematically developed and formalized by British psychologist Philip Nicholas Johnson-Laird in the 1980s through his Mental Models Theory. In his influential work ”Mental Models: Towards a Cognitive Science of Language, Inference, and Consciousness” (1983), Johnson-Laird demonstrated that human reasoning operates through the construction and manipulation of mental models—internal representations that preserve the structural relationships of the situations they represent (johnson1983mental, ). According to this theory, when humans engage in deductive reasoning, inductive inference, or counterfactual thinking, they mentally simulate different scenarios by constructing and testing alternative models of possible worlds.
Johnson-Laird’s framework established several key principles that directly parallel contemporary AI world models: mental models are finite representations of potentially infinite domains, they capture structural relationships rather than superficial details, and they enable predictive simulation of alternative scenarios. These principles have become fundamental to understanding how both human and artificial agents can efficiently represent and reason about complex environments.
2.3. Categorization
@@
  • Applications of world models (Section 5): World models have demonstrated wide-ranging applications across diverse domains, spanning game intelligence, embodied agents, urban systems, and societal modeling. These domains—represented respectively by generative games, robotics, autonomous driving, and social simulacra—illustrate how world models bridge perception, reasoning, and imagination across both virtual and physical environments. We explore how the integration of world models in these domains advances theoretical understanding and practical innovation alike, underscoring their transformative potential in shaping intelligent systems.
@@ 3.1.1. World model in model-based RL
[…]
@@ 3.1.2. World model with language backbone
Recent works have continued to enhance this paradigm by addressing specific challenges, such as those found in web navigation, as exemplified by the research on web agents (chae2024web, ; qiao2024agent, ). Chae et al. (chae2024web, ) developed a World-model-augmented (WMA) agent that improves web navigation by predicting action outcomes via a novel transition-focused observation abstraction, addressing complex issues like irreversible actions such as purchasing non-refundable flight tickets. Similarly, Qiao et al. (qiao2024agent, ) proposed a parametric World Knowledge Model (WKM) that provides agents with both prior global knowledge and dynamic local knowledge, effectively mitigating common problems like blind trial-and-error and hallucinatory actions.
@@
This modular approach has also been successfully applied to specific domains, such as web navigation. Gu et al. (gu2024your, ) proposed WebDreamer, a model-based planning framework that uses a specialized LLM as a world model to simulate actions, achieving competitive performance with significantly higher efficiency than tree-search methods on the web. In a different approach, Tang et al. (tang2024worldcoder, ) introduced WorldCoder, a model-based agent that builds and refines its world model by writing and editing a Python program, demonstrating greater sample and compute efficiency than existing methods.
@@ 3.2. World Knowledge Learned by Models
BLEnD (myung2024blend, ) 2024 NeurIPS Language Benchmark
PIGEON (lan2025open, ) 2025 ACL Findings Language Prediction
LocalGPT (lan2025benchmarking, ) 2025 KDD Language Benchmark
@@
GeoLLM (manvi2023geollm, ) 2024 ICLR Language Understanding
GeoLLM-Bias (manvi2024large, ) 2024 ICML Language Understanding
@@
CityGPT (feng2024citygpt, ) 2025 KDD Language Understanding
CityBench (feng2024citybench, ) 2025 KDD Language&Vision Benchmark
UrbanLLaVA (feng2025urbanllava, ) 2025 ICCV Language&Vision Understanding
GPS-To-Image (feng2025gps, ) 2025 CVPR Vision Generation
Ai’s Blind Spots (beneduce2025ai, ) 2025 arxiv vision Generation
AgentMove (feng2025agentmove, ) 2025 NAACL Language Prediction
GLOBE (li2025recognition, ) 2025 arxiv Language&Vision Understanding
@@
WM-ABench (gao2025vision, ) 2025 ACL Findings Vision Benchmark
Spatial457 (spatiallm, ) 2025 CVPR Vision Benchmark
Thinking in Space (yang2025thinking, ) 2025 CVPR Vision Benchmark
@@
EAI (mozikov2024eai, ) 2024 NeurIPS Language Benchmark
LLM-ToM (kosinski2024evaluating, ) 2024 PNAS Lanauge Benchmark
SafeWorld (yin2024safeworld, ) 2024 NeurIPS Lanuage Benchmark
100 languages (vayani2025all, ) 2025 CVPR Language Benchmark
@@
After pretraining on large-scale web text and books (touvron2023llama, ; chatgpt, ), large language models attain extensive knowledge about the real world and common sense relevant to daily life. This embedded knowledge is considered crucial for their remarkable ability to generalize and perform effectively in real-world tasks. For instance, researchers leverage the common sense of large language models for task planning (zhao2024large, ), robot control (huang2022inner, ), and image understanding (liu2024visual, ). Furthermore, Li et al.(li2024geometryconceptssparseautoencoder, ) discover brain-like structures of world knowledge embedded in the high-dimensional vectors that represent the universe of concepts in large language models. Also, Li et al.(li2024vision, ) demonstrate that language models partially converge towards representations isomorphic to those of vision models. Building upon this extensive knowledge of human daily life (myung2024blend, ), LLMs have been successfully applied in real-world scenarios. For example, by leveraging this prior knowledge to provide semantic information for everyday human activities, they have proven effective in domains such as local life services (lan2025benchmarking, ; lan2025open, ).
@@ 3.2.1. Knowledge of the Global Physical World
While large language models do acquire some implicit knowledge of the real world (gurnee2023language, ; li2024geometryconceptssparseautoencoder, ), the quality of this knowledge remains questionable (roberts2023gpt4geo, ; feng2024citygpt, ). For example, Feng et al. (feng2024citygpt, ; feng2025urbanllava, ) find that the urban knowledge embedded in large language models is often coarse and inaccurate. To address this, they propose an effective framework to improve the acquisition of urban knowledge of specific cities in large language models. Building on the global geospatial knowledge embedded in LLMs, researchers are now applying this prior world knowledge to overcome the generalization challenges faced by previous methods, for instance, AgentMove (feng2025agentmove, ) for global mobility prediction, GPS-to-Image (feng2025gps, ; beneduce2025ai, ) for generating geographically-aware and stylistically-controllable images, and GLOBE for knowledge based image geo-localization (li2025recognition, ).
From a long-term perspective, we can see that although large language models have demonstrated the ability to capture certain aspects of real-world knowledge (gurnee2023language, ; li2024geometryconceptssparseautoencoder, ; roberts2023gpt4geo, ), it is clear that further efforts are needed to enhance this knowledge to enable broader and more reliable real-world applications (feng2025survey, ).
@@ 3.2.2. Knowledge of the Local Physical World
Recent studies explore actively encouraging models to learn abstract knowledge through cognitive map-like processes across various environments. For example, Cornet et al. (gornet2024automated, ) show that in a simplified Minecraft world, visual predictive coding lets an agent build a spatial cognitive map purely from pixels. Once trained, the latent map encodes its metric distance to any target, enabling accurate rollout of future observations. Lin et al. (lin2024learningmodelworldlanguage, ) investigate teaching models to understand the game environments through a world model learning procedure, specifically by predicting the subsequent frame of the environment. In this way, the model can generate better actions in dynamic environments. Moreover, Jin et al. (jinemergent, ) find that language models can learn the emergent representations of program semantics by predicting the next token. Recently, researchers (gao2025vision, ; spatiallm, ; yang2025thinking, ) have extended these studies to more realistic settings, revealing a substantial gap in the ability of large language model-based methods to construct precise models, even for simple local environments.